# Minimum Volume Conformal Sets for Multivariate Regression

### v1

@@ 2.2 Learning $p$-norm prediction sets
Here, the volume of the $p$-norm-based set is still given by $\textrm{Vol}(\mathbb{B}(p,M,\mu))=\lambda(B_{\|\cdot\|_{p}}(1))\cdot\det(M)^{-1}$, where $\lambda(B_{\|\cdot\|_{p}}(1))$ is the Lebesgue measure of the unit $p$-norm ball. Its explicit formula is given by [13],
@@
where $\Gamma(\cdot)$ denotes the gamma function [4]. For brevity, we will retain the notation $\lambda(B_{\|\cdot\|_{p}}(1))$ throughout the remainder of the paper. Building on the reasoning in Proposition 2.1, we derive the following corollary, which provides a more structured and tractable reformulation of the MVCS problem (12).
@@
over $A\in\mathbb{R}^{k\times k}$, $\mu\in\mathbb{R}^{k}$, and $p\in\mathbb{R}$. For this objective, standard first-order optimization algorithms, such as gradient descent, can be applied. For our experiments, we update gradients using the Adam optimizer [24] in PyTorch, with a learning rate scheduler (see Appendix E). After optimization, the final transformation is recovered as $\Lambda^{*}=A^{*}A^{*\top}$.
Despite the challenges induced by the nonconvexity [30], we observe empirically that the joint optimization over $(A,\mu,p)$ with nonconvex gradient descent performs remarkably well in practice. In particular, our numerical experiments demonstrate that the learned values of $p$ often yield well-calibrated prediction sets that align with the intrinsic geometric properties of the data. We illustrate this in Figure 3, where the left panel compares the MVCS obtained via problem (14) to the DCA approach with fixed $p=2$. While both maintain the prescribed coverage level, optimizing $p$ yields a more flexible shape that better fits the data distribution. The right panel further demonstrates this effect in an anisotropic setting, where the learned $p=0.58$ results in an elongated prediction set, adapting to the underlying structure.
@@
where we have used the expression of the volume provided in (16). Now, due the cardinality constraint, for given $R,\mu,\{D_{j}\}_{j=1}^{m}$, and $\{p_{j}\}_{j=1}^{m}$, the optimal $\nu^{\star}(R,\mu,\{D_{j}\}_{j=1}^{m},\{p_{j}\}_{j=1}^{m})$ is equal to the $(n-r+1)$-th smallest value in the set $\left\{d(y_{i},\mu;{R,\{D_{j}\}_{j=1}^{m}})\right\}_{i=1}^{n}$, or, equivalently, the $r$-th largest value, i.e., $\nu^{\star}=\sigma_{r}\left\{d(y_{i},\mu;{R,\{D_{j}\}_{j=1}^{m}})\right\}$. Plugging this into (18), we obtain problem (17) in the statement of the proposition. ∎
@@
To enable unconstrained optimization over the rotation matrix $R$, we introduce a parameterization based on the QR decomposition. Specifically, we optimize over a full matrix $Q\in\mathbb{R}^{k\times k}$ and extract the rotation matrix via the decomposition $Q=R_{1}R_{2}$, where $R_{1}$ is an orthogonal matrix and $R_{2}$ is upper triangular. Since $R_{1}$ may have determinant $-1$, we ensure that $R$ remains a proper rotation matrix in $\mathrm{SO}(k)$ by setting $R=R_{1}$ if $\det R_{1}=1$, and $R=I_{-}R_{1}$ if $\det R_{1}=-1$, where $I_{-}$ is a diagonal matrix with all ones except for a $-1$ in the first diagonal entry.
@@ 3 Local Adaptivity in Multivariate Regression
where we have used the expression of the volume provided in (4). Now, due the cardinality constraint, for given $\Lambda,p,\theta$, the optimal $\nu^{\star}(\Lambda,p,\theta)$ is equal to the $(n-r+1)$-th smallest value in the set $\left\{\|\Lambda(x_{i})(y_{i}-f_{\theta}(x_{i}))\|_{p}\right\}_{i=1}^{n}$, or, equivalently, the $r$-th largest value, i.e., $\nu^{\star}=\sigma_{r}\left\{\|\Lambda(x_{i})(y_{i}-f_{\theta}(x_{i}))\|_{p}\right\}$. Plugging this into (23), we obtain problem (22). ∎
The first term in the objective function ensures that the average volume of the prediction sets across all feature values is minimized, rather than optimizing a single global shape. This accounts for varying uncertainty levels throughout the feature space. The second term enforces the coverage constraint by selecting the $r$-th largest nonconformity score, ensuring that at least $n-r+1$ points remain within their respective sets.
@@ 4 Conformalized Minimum-Volume Prediction Sets
The conformalization procedure we adopt follows the standard split conformal prediction framework [59, 40], with the key difference being our choice of nonconformity scores, which are derived from the locally adaptive MVCS sets presented in Section 3. We consider a setting where we are given a dataset of $n$ i.i.d. samples $(X_{i},Y_{i})\sim\mathbb{P}$, where $\mathbb{P}$ is the unknown joint distribution of covariates $X_{i}\in\mathbb{R}^{d}$ and responses $Y_{i}\in\mathbb{R}^{k}$. Our goal is to construct covariate-dependent prediction sets $C(x)$ that satisfy the coverage guarantee
@@ 5.2 Comparative baselines
We compare our method (single-norm MVCS) against three established baselines: Naïve Quantile Regression [10] (naïve QR), the empirical covariance matrix [23] (emp. cov.), and the local empirical covariance method [35] (loc. emp. cov.). Below, we summarize these approaches.
  • Naïve quantile regression. This strategy involves performing quantile regression using the pinball loss [50] for the quantiles ${\tilde{\alpha}}/{2}$ and $1-{\tilde{\alpha}}/{2}$ along all axes on a training dataset. To avoid overcovering, we select $\tilde{\alpha}=2(1-(1-\alpha)^{1/k})$, as suggested by previous work [10] to fit the quantile networks $q$. The prediction regions are then conformalized on a calibration dataset using the score
@@
  • Empirical covariance matrix. A widely used alternative constructs ellipsoidal prediction regions by leveraging the empirical covariance structure of the residuals [23]. The estimated covariance matrix $\hat{\Sigma}$ is computed from the training residuals, and the prediction region is shaped using the Mahalanobis norm. The nonconformity score is given by
@@
  • Local empirical covariance matrix. While the global covariance structure captures overall variability, it fails to adapt to local heteroskedasticity. To address this limitation, a local empirical covariance approach [35] estimates $\hat{\Sigma}_{x}$ using only the $m$ nearest neighbors of $x$ in the training dataset. The nonconformity score is then computed as
@@ 5.3 Synthetic dataset
A key consideration in conformal prediction is whether a given strategy maintains conditional coverage. While marginal coverage guarantees ensure that $\mathrm{Prob}\left\{Y\in C(X)\right\}\geq 1-\alpha$, conditional validity (i.e., requiring $\mathrm{Prob}\{Y\in C(X)\mid X\}=1-\alpha$) is generally unattainable without strong assumptions on the data distribution [15]. Aggressively minimizing volume can degrade conditional validity [21], as our optimization objective (22) does not explicitly enforce constraints on the $r$-largest values. However, the neural network parameterization promotes generalization, leading to empirically favorable results despite the lack of theoretical convergence guarantees due to nonconvexity.
@@
Figure 8: (Left) Conformal prediction sets for a target coverage level of $0.99$. (Right) Conformal prediction sets for a target coverage level of $0.90$, with $43\alpha/$ 4 of the data consisting of outliers.
@@
Our loss function is inherently robust to outliers, up to a fraction of $\alpha$. To illustrate this, we replicate the previous experiment while introducing extreme values for a fraction $43\alpha/$ 4 of the data points. [^3] These extreme values are fixed at 10. We compare our approach with the pinball loss and the standard nonconformity score, using the same evaluation framework. The pinball loss is trained on the quantiles $\alpha/2$ and $21-\alpha/$ 2, but under contamination, its optimal quantiles shift, leading to unreliable prediction sets. In particular, when the proportion of outliers in the upper quantile exceeds $\alpha/2$, the estimated $(1-\alpha)$-quantile is dominated by extreme values, compromising the validity of the prediction sets. This highlights the limitations of quantile-based methods in the presence of outliers, whereas our approach remains robust, maintaining stable and reliable uncertainty quantification.
@@ 5.4 Real datasets
We evaluate our approach on nine benchmark datasets commonly used in multivariate regression studies [10, 14, 60]. Details regarding these datasets, including the response dimensionality and hyperparameter choices, are provided in Appendix D. We compare different strategies for both $\alpha=0.1$ and $\alpha=0.01$. Each dataset is split into four subsets—training, validation, calibration, and test—using respective proportions of 70%, 10%, 10%, and 10%. [^4] Except for the energy dataset, for which the data splitting is 55%, 15%, 15%, 15% due to the smaller amount of samples. To ensure robust evaluation, results are averaged over 10 independent runs, discarding the highest and lowest volume values. Additionally, each volume is scaled by taking its $1/k$-th power to normalize for dimensionality.
To preprocess the data, we apply a quantile transformation to both the covariates ($X$) and responses ($Y$) using scikit-learn [41], ensuring normalization across datasets. Our approach consistently minimizes the volume of prediction sets while maintaining valid marginal coverage. Specifically, we achieve the best average volume in sixteen out of eighteen experiments while preserving the desired coverage level (Tables 4 and 6).
@@
Bias correction $99.1\pm 0.2$ $99.3\pm 0.2$ $99.1\pm 0.4$ $99.4\pm 0.4$
CASP $99.2\pm 0.1$ $99.0\pm 0.1$ $99.0\pm 0.1$ $99.1\pm 0.1$
Energy $99.8\pm 0.4$ $99.8\pm 0.4$ $99.8\pm 0.4$ $99.4\pm 0.8$
House $99.1\pm 0.2$ $99.0\pm 0.3$ $99.1\pm 0.2$ $99.0\pm 0.2$
rf1 $99.0\pm 0.3$ $99.0\pm 0.3$ $99.0\pm 0.5$ $99.0\pm 0.3$
rf2 $99.0\pm 0.4$ $99.0\pm 0.5$ $99.1\pm 0.1$ $99.2\pm 0.3$
scm1d $99.2\pm 0.3$ $99.3\pm 0.3$ $99.3\pm 0.3$ $99.3\pm 0.3$
scm20d $99.1\pm 0.2$ $99.3\pm 0.2$ $99.3\pm 0.2$ $99.1\pm 0.3$
@@ A Proof of convexity in the DC formulation
To establish the convexity of $f(\Lambda,\mu)$, we analyze its two components. The first term, $-\log\det\Lambda$, is a standard convex function over the space of positive semidefinite matrices. The second term, $r\overline{\sigma}_{r}\{\|\Lambda y_{i}+\eta\|\}$, can be shown to be convex by expressing it as the following optimization problem
@@ B Synthetic dataset generation
$\tilde{w}_{i}(X)=\frac{1}{|X-X_{i}|^{4}},\quad w_{i}(X)=\frac{\tilde{w}_{i}(X)}{\sum_{j}\tilde{w}_{j}(X)}.$

### v2

@@ 2.2 Learning $p$-norm prediction sets
Here, the volume of the $p$-norm-based set is still given by $\textrm{Vol}(\mathbb{B}(p,M,\mu))=\lambda(B_{\|\cdot\|_{p}}(1))\cdot\det(M)^{-1}$, where $\lambda(B_{\|\cdot\|_{p}}(1))$ is the Lebesgue measure of the unit $p$-norm ball. Its explicit formula is given by Dirichlet [1839],
@@
where $\Gamma(\cdot)$ denotes the gamma function Artin [2015]. For brevity, we will retain the notation $\lambda(B_{\|\cdot\|_{p}}(1))$ throughout the remainder of the paper. Building on the reasoning in Proposition 2.1, we derive the following corollary, which provides a more structured and tractable reformulation of the MVCS problem (12).
@@
over $A\in\mathbb{R}^{k\times k}$, $\mu\in\mathbb{R}^{k}$, and $p\in\mathbb{R}$. For this objective, standard first-order optimization algorithms, such as gradient descent, can be applied. For our experiments, we update gradients using the Adam optimizer Kingma and Ba [2014] in PyTorch, with a learning rate scheduler (see Appendix E). After optimization, the final transformation is recovered as $\Lambda^{*}=A^{*}A^{*\top}$.
Despite the challenges induced by the nonconvexity Lee et al. [2016], we observe empirically that the joint optimization over $(A,\mu,p)$ with nonconvex gradient descent performs remarkably well in practice. In particular, our numerical experiments demonstrate that the learned values of $p$ often yield well-calibrated prediction sets that align with the intrinsic geometric properties of the data. We illustrate this in Figure 3, where the left panel compares the MVCS obtained via problem (14) to the DCA approach with fixed $p=2$. While both maintain the prescribed coverage level, optimizing $p$ yields a more flexible shape that better fits the data distribution. The right panel further demonstrates this effect in an anisotropic setting, where the learned $p=0.58$ results in an elongated prediction set, adapting to the underlying structure.
@@
where we have used the expression of the volume provided in (16). Now, due the cardinality constraint, for given $R,\mu,\{D_{j}\}_{j=1}^{m}$, and $\{p_{j}\}_{j=1}^{m}$, the optimal $\nu^{\star}(R,\mu,\{D_{j}\}_{j=1}^{m},\{p_{j}\}_{j=1}^{m})$ is equal to the (nr+1)(n-r+1)-th smallest value in the set $\left\{d(y_{i},\mu;{R,\{D_{j}\}_{j=1}^{m}})\right\}_{i=1}^{n}$, or, equivalently, the $r$-th largest value, i.e., $\nu^{\star}=\sigma_{r}\left\{d(y_{i},\mu;{R,\{D_{j}\}_{j=1}^{m}})\right\}$. Plugging this into (18), we obtain problem (17) in the statement of the proposition. ∎
@@
To enable unconstrained optimization over the rotation matrix $R$, we introduce a parameterization based on the QR decomposition. Specifically, we optimize over a full matrix $Q\in\mathbb{R}^{k\times k}$ and extract the rotation matrix via the decomposition $Q=R_{1}R_{2}$, where $R_{1}$ is an orthogonal matrix and $R_{2}$ is upper triangular. Since $R_{1}$ may have determinant 1-1, we ensure that $R$ remains a proper rotation matrix in $\mathrm{SO}(k)$ by setting $R=R_{1}$ if $\det R_{1}=1$, and $R=I_{-}R_{1}$ if $\det R_{1}=-1$, where $I_{-}$ is a diagonal matrix with all ones except for a 1-1 in the first diagonal entry.
@@ 3 Local Adaptivity in Multivariate Regression
where we have used the expression of the volume provided in (4). Now, due the cardinality constraint, for given $\Lambda,p,\theta$, the optimal $\nu^{\star}(\Lambda,p,\theta)$ is equal to the (nr+1)(n-r+1)-th smallest value in the set $\left\{\|\Lambda(x_{i})(y_{i}-f_{\theta}(x_{i}))\|_{p}\right\}_{i=1}^{n}$, or, equivalently, the $r$-th largest value, i.e., $\nu^{\star}=\sigma_{r}\left\{\|\Lambda(x_{i})(y_{i}-f_{\theta}(x_{i}))\|_{p}\right\}$. Plugging this into (23), we obtain problem (22). ∎
The first term in the objective function ensures that the average volume of the prediction sets across all feature values is minimized, rather than optimizing a single global shape. This accounts for varying uncertainty levels throughout the feature space. The second term enforces the coverage constraint by selecting the $r$-th largest nonconformity score, ensuring that at least nr+1n-r+1 points remain within their respective sets.
@@ 4 Conformalized Minimum-Volume Prediction Sets
The conformalization procedure we adopt follows the standard split conformal prediction framework Vovk et al. [2005]; Papadopoulos et al. [2002], with the key difference being our choice of nonconformity scores, which are derived from the locally adaptive MVCS sets presented in Section 3. We consider a setting where we are given a dataset of $n$ i.i.d. samples $(X_{i},Y_{i})\sim\mathbb{P}$, where $\mathbb{P}$ is the unknown joint distribution of covariates $X_{i}\in\mathbb{R}^{d}$ and responses $Y_{i}\in\mathbb{R}^{k}$. Our goal is to construct covariate-dependent prediction sets $C(x)$ that satisfy the coverage guarantee
@@ 5.2 Comparative baselines
We compare our method (single-norm MVCS) against three established baselines: Naïve Quantile Regression Dheur et al. [2025] (naïve QR), the empirical covariance matrix Johnstone and Cox [2021] (emp. cov.), and the local empirical covariance method Messoudi et al. [2022] (loc. emp. cov.). Below, we summarize these approaches.
  • Naïve quantile regression. This strategy involves performing quantile regression using the pinball loss Steinwart and Christmann [2011] for the quantiles ${\tilde{\alpha}}/{2}$ and $1-{\tilde{\alpha}}/{2}$ along all axes on a training dataset. To avoid overcovering, we select $\tilde{\alpha}=2(1-(1-\alpha)^{1/k})$, as suggested by previous work Dheur et al. [2025] to fit the quantile networks $q$. The prediction regions are then conformalized on a calibration dataset using the score
@@
  • Empirical covariance matrix. A widely used alternative constructs ellipsoidal prediction regions by leveraging the empirical covariance structure of the residuals Johnstone and Cox [2021]. The estimated covariance matrix $\hat{\Sigma}$ is computed from the training residuals, and the prediction region is shaped using the Mahalanobis norm. The nonconformity score is given by
@@
  • Local empirical covariance matrix. While the global covariance structure captures overall variability, it fails to adapt to local heteroskedasticity. To address this limitation, a local empirical covariance approach Messoudi et al. [2022] estimates $\hat{\Sigma}_{x}$ using only the $m$ nearest neighbors of $x$ in the training dataset. The nonconformity score is then computed as
@@ 5.3 Synthetic dataset
A key consideration in conformal prediction is whether a given strategy maintains conditional coverage. While marginal coverage guarantees ensure that $\mathrm{Prob}\left\{Y\in C(X)\right\}\geq 1-\alpha$, conditional validity (i.e., requiring $\mathrm{Prob}\{Y\in C(X)\mid X\}=1-\alpha$) is generally unattainable without strong assumptions on the data distribution Foygel Barber et al. [2021]. Aggressively minimizing volume can degrade conditional validity Huang et al. [2023], as our optimization objective (22) does not explicitly enforce constraints on the $r$-largest values. However, the neural network parameterization promotes generalization, leading to empirically favorable results despite the lack of theoretical convergence guarantees due to nonconvexity.
@@
Figure 8: (Left) Conformal prediction sets for a target coverage level of $0.99$. (Right) Conformal prediction sets for a target coverage level of $0.90$, with $3\alpha/4$ of the data consisting of outliers.
@@
Our loss function is inherently robust to outliers, up to a fraction of $\alpha$. To illustrate this, we replicate the previous experiment while introducing extreme values for a fraction $3\alpha/4$ of the data points. [^3] These extreme values are fixed at 10. We compare our approach with the pinball loss and the standard nonconformity score, using the same evaluation framework. The pinball loss is trained on the quantiles $\alpha/2$ and $1-\alpha/2$, but under contamination, its optimal quantiles shift, leading to unreliable prediction sets. In particular, when the proportion of outliers in the upper quantile exceeds $\alpha/2$, the estimated $(1-\alpha)$-quantile is dominated by extreme values, compromising the validity of the prediction sets. This highlights the limitations of quantile-based methods in the presence of outliers, whereas our approach remains robust, maintaining stable and reliable uncertainty quantification.
@@ 5.4 Real datasets
We evaluate our approach on nine benchmark datasets commonly used in multivariate regression studies Dheur et al. [2025]; Feldman et al. [2023]; Wang et al. [2023]. Details regarding these datasets, including the response dimensionality and hyperparameter choices, are provided in Appendix D. We compare different strategies for both $\alpha=0.1$ and $\alpha=0.01$. Each dataset is split into four subsets—training, validation, calibration, and test—using respective proportions of 70%, 10%, 10%, and 10%. [^4] Except for the energy dataset, for which the data splitting is 55%, 15%, 15%, 15% due to the smaller amount of samples. To ensure robust evaluation, results are averaged over 10 independent runs, discarding the highest and lowest volume values. Additionally, each volume is scaled by taking its $1/k$-th power to normalize for dimensionality.
To preprocess the data, we apply a quantile transformation to both the covariates ($X$) and responses ($Y$) using scikit-learn Pedregosa et al. [2011], ensuring normalization across datasets. Our approach consistently minimizes the volume of prediction sets while maintaining valid marginal coverage. Specifically, we achieve the best average volume in sixteen out of eighteen experiments while preserving the desired coverage level (Tables 4 and 6).
@@
Bias correction $9.1\pm 0.2 99$$3\pm 0.2 99.$$1\pm 0.4 99.$$9.4\pm 0.4$
CASP $9.2\pm 0.1 99$$\pm 0.1 99.0±0$$\pm 0.1 99$$9.1\pm 0.1$
Energy $9.8\pm 0.4 99$$8\pm 0.4 99.$$8\pm 0.4 99.$$9.4\pm 0.8$
House $9.1\pm 0.2 99$$0\pm 0.3 99.$$1\pm 0.2 99.$$9.0\pm 0.2$
rf1 $9.0\pm 0.3 99$$\pm 0.3 99.0±0$$\pm 0.5 99$$9.0\pm 0.3$
rf2 $9.0\pm 0.4 99$$0\pm 0.5 99.$$1\pm 0.1 99.$$9.2\pm 0.3$
scm1d $9.2\pm 0.3 99$$3\pm 0.3 99.$$3\pm 0.3 99.$$9.3\pm 0.3$
scm20d $9.1\pm 0.2 99$$3\pm 0.2 99.$$3\pm 0.2 99.$$9.1\pm 0.3$
@@ A Proof of convexity in the DC formulation
To establish the convexity of $f(\Lambda,\mu)$, we analyze its two components. The first term, $\log\det\Lambda$, is a standard convex function over the space of positive semidefinite matrices. The second term, $r\overline{\sigma}_{r}\{\|\Lambda y_{i}+\eta\|\}$, can be shown to be convex by expressing it as the following optimization problem
@@ B Synthetic dataset generation
$\tilde{w}_{i}(X)=\frac{1}{\|X-X_{i}\|^{4}},\quad w_{i}(X)=\frac{\tilde{w}_{i}(X)}{\sum_{j}\tilde{w}_{j}(X)}.$